You must have both the minutes and the seconds correct for My Mathab to count your answer as correct you calculate a whole number of mes, such as 2 minutes, you will need 2 min 0 sec for My Mathlab to
give you credit

The doctor orders morphine sultate 10 mg IVP stat the recommended rate of administration is 15 mg over 4-5 minutes Calculate the appropriate length of time for the ordered dose. Convert your answers to minutes and seconds.

Find the maximum length of time


__?___Minutes __?___seconds is the maximum amount of time
Answer #1

240 minutes and 100seconds is the maximum amount of time
recommended rate is 15mg over 4-5minutes
Doctor order 10mg
so 15mg/4-5
1hour =60minutes
4-5 hours= 4X 60=240minutes
10 X 15 (standard drip factor)/240 =10 X15/240 =0.6drops /minute
1 minute =60 seconds
60 sec divided by 0.6 = 100seconds.